OpenStudy (anonymous):

Which of the following is not a way to represent the solution of the inequality 2(x - 2) less than or greater to 2? A number line with a closed circle on 3 and shading to the right. A number line with a closed circle on 3 and shading to the left. 3 greater than or equal to x x less than or greater to 3

OpenStudy (anonymous):

@Mertsj @hartnn

hartnn (hartnn):

"less than or greater " ??? what does that mean ?

OpenStudy (anonymous):

Sorry, it's a typo. It's actually 'Less than or equal to"

hartnn (hartnn):

\(\large 2(x-2) \le 2\) can you isolate 'x' from here ?

hartnn (hartnn):

you can start by dividing 2 on both sides.

OpenStudy (anonymous):

I don't know how to do that type.

OpenStudy (anonymous):

Oh, okay

OpenStudy (anonymous):

Wait, @hartnn How would I divide that?

hartnn (hartnn):

\(\large 2(x-2) \le 2 \implies \large (x-2) \le 1\)

hartnn (hartnn):

now just add 2 on both sides.

OpenStudy (anonymous):

x-4 <= 3?

hartnn (hartnn):

if you add 2 to x-2 you get x-2+2 =x only

OpenStudy (anonymous):

oh okay

hartnn (hartnn):

so, x<= 3 means which option ?

hartnn (hartnn):

i mean x<=3 is not represented by which option ?

OpenStudy (anonymous):

The first one! :)

hartnn (hartnn):

correct ! :D